# Machine Learning Symmetry Discovery for Classical Mechanics

### v1

@@ Abstract
In this study, we propose a data-driven, deep-learning-based Machine-Learning Symmetry Discovery (MLSD) algorithm to automate the discovery of continuous Lie group symmetries in classical mechanical systems from their time-evolution trajectory data. MLSD uses neural networks (NNs) to predict conserved physical quantities that implement symmetry transformations of the phase space coordinates. After training, MLSD is able to identify the Lie algebra, particularly non-abelian ones, as indicated by the Lie algebra structure coefficients. To demonstrate the effectiveness of the MLSD method, we applied it to simulated data from the classical three-dimensional Kepler problem and the harmonic oscillator. The results show that the algorithm successfully identified the hidden symmetry groups of both systems.
@@ I Introduction
While the discovery of symmetries is invaluable, it is often a challenging task that typically requires the sophisticated expertise of physicists. To address this challenge, advancing field of machine learning[5, 6, 7] offers a promising solution for facilitating the comprehensive discovery of symmetries in complex systems. One significant advantage of machine learning in this context is its ability to learn and compress from large volumes of data. This data-driven approach can sometimes surpass human intuition, as machine learning algorithms can identify patterns and correlations that may not be manifest to human researchers. By harnessing the vast amounts of data generated in scientific research, machine learning holds the potential to accelerate discovery and provide new insights into the fundamental laws and symmetries governing physical systems[8, 9, 10].
Previous work in this area has achieved significant success in uncovering continuous symmetries in classical mechanical systems[11, 12, 13, 14, 15, 16, 17]. However, a more challenging extension is to develop a systematic framework capable of automatically discovering possible continuous Lie group symmetries, particularly non-Abelian ones, from a dataset, rather than relying on an analytical form of the Hamiltonian. In this study, we propose a Machine-Learning Symmetry Discovery (MLSD) algorithm designed to explore all possible continuous Lie group symmetry transformations in classical mechanical systems using their time-evolution datasets. The MLSD algorithm identifies conservation laws that correspond to continuous symmetry transformations. After training, MLSD outputs a three-way tensor $f$, representing the Lie algebra structure coefficients of the discovered set of symmetry transformations. The generality of these structure coefficients enables MLSD to identify both Abelian and non-Abelian symmetries.
@@ II Methodology
Our task of symmetry discovery differs from those explored in previous studies. For example, [11, 14] focuses on identifying transformations that preserve time evolution trajectories, while [12, 13] examines whether specific proposed transformations can be learned by neural networks. In contrast, our approach seeks to uncover symmetry transformations that extend beyond preserving the shape of trajectories and instead focus solely on preserving energy, as similarly discussed in [15, 16, 17], without relying on prior human knowledge. In this section, we begin by reviewing the concept of symmetry in classical systems as it pertains to our study. We then introduce the architecture of the Machine Learning Symmetry Discovery (MLSD) framework, which is designed to identify these symmetries.
@@ II.1 Continuous symmetry and symmetry group of a physical system
The time evolution dynamics of a physical system is encoded in the Hamiltonian $H({\bm{x}})$, which depends on the canonical coordinates ${\bm{x}}:=({\bm{q}},{\bm{p}})$, with ${\bm{q}}$ and ${\bm{p}}$ being the conjugate position and momentum variables. A continuous symmetry of the system is defined as the conservation of energy under continuous symmetry transformations generated by a conserved quantity. This can be expressed through the vanishing Poisson bracket commutator between $H({\bm{x}})$ and the conserved quantity $G_{i}({\bm{x}})$, where $i\in\{1,2,...,n\}$ and $n$ denote the dimension of the symmetry group. The identification of the continuous symmetry group (Lie group) follows directly from the Lie algebra structure coefficients $f^{ijk}$, which are obtained from the commutators between the quantities.
@@
The goal of discovering all possible continuous symmetries is equivalent to finding all independent quantity functions $G_{i}({\bm{x}}),\forall i\in\{0,1,2,...,n\}$, along with the structure coefficient $f^{ijk}$ that define the symmetry group. The special $G_{0}$ component corresponds to the Hamiltonian $H$, relating to energy conservation. Its dependence $H({\bm{x}})$ on ${\bm{x}}$ can be learned from a series of time evolution data ${\bm{x}}(t)$, assuming the underlying principle of Hamiltonian dynamics of classical mechanics.
@@ II.2 Learning continuous symmetry transformations
Considering a classical system in spatial dimension $d$, let ${\bm{X}}=\{{\bm{x}}(t)\in\mathbb{R}^{2d}:\dot{{\bm{x}}}=\{{\bm{x}},H\}\}$ be a dataset of time series samples, where each sample is a trajectory of the canonical coordinate ${\bm{x}}(t)$ following the underlying Hamiltonian dynamics driven by $H$. The data can be collected from observation, and the Hamiltonian $H$ may not be known to us. The goal is to uncover the continuous Lie group symmetries of such a classical dynamical system as defined by Eq. (5). To achieve this, we propose the Machine-Learning Symmetry Discovery (MLSD) algorithm, which uses neural networks (NNs) to parametrize each of the conserved quantities in Eq. (5) as $G_{i,\theta}$, where $\theta$ represents the learnable parameters in neural networks.
@@
where $\epsilon$ denotes an infinitesimal variation and $\mathbf{J}$ is the symplectic metric defined in Eq. (4). Based on this principle, we can parametrize a conserved quantity $G_{i,\theta}$ and use its gradient, $\nabla_{{\bm{x}}}G_{i,\theta}$, to generate a symmetry transformation $\delta{\bm{x}}=\mathbf{J}(\nabla_{\bm{x}}G_{i})\epsilon$. In contrast to prior studies that directly model the symmetry transformation $\delta{\bm{x}}$ by neural networks[11, 12, 13, 14, 16], our approach of modeling $\delta{\bm{x}}$ indirectly through gradients of scalar functions $G_{i,\theta}$ is simpler, while inherently respecting the curl-free constraint $\nabla_{{\bm{x}}}\times(\nabla_{{\bm{x}}}G_{i,\theta})=0$ that should otherwise be imposed on $\delta{\bm{x}}$ as well.
@@
    The symmetry discovery process begins by inputting a conjectured dimension $n$ of the Lie group. Therefore, $n+1$ individual neural networks are used to predict each conserved quantity. The structure coefficient $f^{ijk}$ is parametrized to be anti-symmetric, defined as $f^{ijk}=\theta^{ijk}-\theta^{ikj}+\theta^{jki}-\theta^{jik}+\theta^{kij}-\theta^{kji}$, for all $i,j,k\neq 0,$ with the parameter $\theta$ being optimizable. In order to recognize the symmetry group formed by all $G_{i,\theta}$, we train on the dataset ${\bm{X}}$ with the following mean square loss:
@@
    It should be understood that the integrand is evaluated at instantaneously at each time $t$ and then integrated over the trajectory ${\bm{x}}(t)$ sampled from the dataset ${\bm{X}}$.
In addition to $\mathcal{L}_{\text{H}}$ and $\mathcal{L}_{\text{G}}$, we introduce a third term into the loss function to maximize the linear independence of the set of vectors $\nabla_{{\bm{x}}}G_{i,\theta}:i\in\{1,2,...,n\}$. Otherwise, the algorithm might converge to trival solutions where several $G_{i,\theta}$ correspond to equivalent conserved quantities, which we should try to avoid. For this purpose, we introduce the following $n\times 2d$ matrix
$\mathbf{M}({\bm{x}})=(\mathbf{J}(\widehat{\nabla_{{\bm{x}}}G_{0,\theta}}),\widehat{\nabla_{{\bm{x}}}G_{1,\theta}},...,\widehat{\nabla_{{\bm{x}}}G_{n,\theta}})^{\intercal},$ (9)
where each $\widehat{\nabla_{{\bm{x}}}G_{i,\theta}}=\nabla_{{\bm{x}}}G_{i,\theta}/{\lVert\nabla_{{\bm{x}}}G_{i,\theta}\rVert}$ denotes the normalized column vector. The linear independence of the learned transformations in Eq. (6) is equivalent to non-vanishing product of the top $\min(2d,n)$ [^1] If $12d\leq n+$ 1, then finite value of $\mathcal{L}_{\text{I}}$ indicates all the vectors in $\mathbf{M}({\bm{x}})$ have fully spanned the $2d$ dimension space; otherwise, the $n+1$ vectors in $\mathbf{M}({\bm{x}})$ are linearly independent to each other. eigenvalues of $\mathbf{M}({\bm{x}})\mathbf{M}^{\intercal}({\bm{x}})$. Denoting the eigenvalues of matrix $\mathbf{M}({\bm{x}})\mathbf{M}^{\intercal}({\bm{x}})$ as $\lambda^{i}({\bm{x}})$, the third term is written as:
$\mathcal{L}_{\text{I}}=-\mathop{\mathbb{E}}_{{\bm{x}}(t)\in{\bm{X}}}\int\frac{1}{N}\text{log}\Big{(}\prod_{i=1}^{\min(2d,n)}\lambda^{i}({\bm{x}})\Big{)}\mathrm{d}t.$ (10)
Since all the vectors defined in matrix $\mathbf{M}({\bm{x}})$ have been normalized, the condition $\prod_{i=1}^{\min(2d,n)}\lambda^{i}({\bm{x}})>0$ indicates maximal independence of the learned symmetry transformations. Consequently, in practice, the independence term in Eq. (10) is scaled by a very small factor when added to the overall loss function to prevent blowing up the total loss function.
In summary, the overall loss function is given by:
@@
with the factor $\beta\ll\alpha\simeq 1$.
After training, if the given dataset ${\bm{X}}$ corresponds to a system with continuous $n$ dimensional symmetry, the first two loss terms, Eq. (7) and Eq. (8), should converge to zero, while the third term, Eq. (10), remains finite. Conversely, if the system lacks an $n$ dimensional symmetry, the first two loss terms will remain finite regardless of how small $\beta$ is set. The specific type of symmetry group can be identified by extracting the structure coefficients $f^{ijk}$ , as detailed in Appendix B.
@@
[…]
@@ III Empirical results
In this section, we aim to implement MLSD to discover symmetries in the harmonic oscillator and Kepler problem, summarized in Tab. 1. A detailed analysis of $\mathrm{SU}(3)$ symmetry of harmonic oscillator and $\mathrm{SO}(4)$ symmetry of Kepler problem is provided in the Appendix A. We demonstrate that MLSD can successfully identify the hidden symmetries in both systems. Additionally, in the harmonic oscillator task, we use quadratic parametrization to show that MLSD can explicitly reconstruct the 8 Gell-Mann matrices that form the $\mathfrak{su}(3)$ Lie algebra. Code is available at [19].

### v2

# Machine Learning Symmetry Discovery for Integrable Hamiltonian Dynamics

@@ Abstract
We propose a data-driven Machine-Learning Symmetry Discovery (MLSD) framework for identifying continuous symmetry generators and their Lie-algebraic structure directly from phase-space trajectory data expressed in canonical coordinates. MLSD parameterizes candidate conserved quantities with neural networks and learns antisymmetric structure coefficients by enforcing Poisson-bracket closure, supplemented by a weak independence regularizer. We validate MLSD on two integrable benchmark systems—the three-dimensional Kepler problem and the three-dimensional isotropic harmonic oscillator—recovering the expected non-Abelian algebras (respectively $\mathfrak{so}(4)$ and $\mathfrak{su}(3)$) up to basis transformations. This work focuses on integrable benchmark dynamics, where global conserved quantities are well-defined and admit compact representations learnable from canonical-coordinate trajectories. Extending symmetry discovery to mixed or chaotic phase-space regimes is an important direction for future work.
@@ I Introduction
While the discovery of symmetries is invaluable, it is often a challenging task that typically requires the sophisticated expertise of physicists. To address this challenge, advancing field of machine learning[11, 1, 3] offers a promising set of tools for assisting symmetry discovery from data in controlled physical settings. One significant advantage of machine learning in this context is its ability to learn and compress from large volumes of data. This data-driven approach provides a systematic way to extract patterns and correlations from data that may not be evident from standard theoretical analysis. By harnessing the vast amounts of data generated in scientific research, machine learning holds the potential to accelerate discovery and provide new insights into the fundamental laws and symmetries governing physical systems[20, 16, 9].
Previous work has demonstrated learning-based approaches to uncovering continuous symmetries in classical mechanical systems, primarily in settings with regular dynamics where conserved quantities admit stable functional representations[22, 14, 13, 23, 17, 4, 19, 12]. In this paper we similarly focus on benchmark Hamiltonian systems with integrable dynamics, where the relevant conserved quantities are well-defined on the sampled phase-space region and admit compact representations. A natural next step within this line of work is to develop a systematic framework for discovering continuous Lie group symmetries, particularly non-Abelian ones, directly from trajectory data, without assuming prior analytic knowledge of the conserved quantities.
In this study, we propose a Machine-Learning Symmetry Discovery (MLSD) algorithm designed to identify continuous Lie group symmetries in classical mechanical systems with integrable Hamiltonian dynamics, using simulated time-evolution trajectory data. The MLSD algorithm identifies conservation laws that correspond to continuous symmetry transformations. After training, MLSD outputs a three-way tensor $f$, representing the Lie algebra structure coefficients of the discovered set of symmetry transformations. The structure coefficients provide a basis-independent representation that allows MLSD to identify both Abelian and non-Abelian symmetry algebras within the considered setting.
Within this deliberately scoped benchmark setting, MLSD provides an end-to-end pipeline that (i) learns candidate conserved quantities from trajectories, (ii) recovers a basis-independent Lie-algebra signature via structure coefficients and Killing-form analysis, and (iii) reports stability and reproducibility diagnostics suitable for systematic comparison across symmetry dimensions.
@@ II Methodology
Our goal of symmetry discovery differs from those explored in previous studies. For example, [22, 23] focuses on identifying transformations that preserve time evolution trajectories, while [14, 13] examines whether specific proposed transformations can be learned by neural networks. In contrast, our approach seeks to uncover symmetry transformations that extend beyond preserving the shape of trajectories and instead focus solely on preserving energy, as similarly discussed in [17, 4, 19], without assuming analytic forms for the conserved quantities. In this section, we begin by reviewing the concept of symmetry in classical systems as it pertains to our study. We then introduce the architecture of the Machine Learning Symmetry Discovery (MLSD) framework, which is designed to identify these symmetries.
@@ II.1 Continuous symmetry and symmetry group of a physical system
The time evolution dynamics of a physical system is encoded in the Hamiltonian $H({\bm{x}})$, which depends on the canonical coordinates ${\bm{x}}:=({\bm{q}},{\bm{p}})$, with $\bm{q}$} and $\bm{p}$} being the conjugate position and momentum variables. A continuous symmetry of the system is defined by the invariance of the Hamiltonian under a continuous family of canonical transformations generated by a conserved quantity. This can be expressed through the vanishing Poisson bracket between $H({\bm{x}})$ and the conserved quantity $G_{i}({\bm{x}})$, where $i\in\{1,2,...,n\}$ and $n$ denote the dimension of the symmetry group. The identification of the continuous symmetry group (Lie group) follows directly from the Lie algebra structure coefficients $f^{ijk}$, which are obtained from the commutators between the quantities.
@@
The goal in this work is to discover a set of independent conserved quantities $G_{i}({\bm{x}})$ and associated structure coefficients $f^{ijk}$ that describe the symmetry algebra, within the expressivity of the chosen parametrization and the phase-space coverage of the dataset. The special $G_{0}$ component corresponds to the Hamiltonian $H$, relating to energy conservation. Its dependence $H({\bm{x}})$ on $\bm{x}$} can be learned from a series of time evolution data ${\bm{x}}(t)$, assuming the underlying principle of Hamiltonian dynamics of classical mechanics.
@@ II.2 Learning continuous symmetry transformations
Considering a classical system in spatial dimension $d$, let ${\bm{X}}=\{{\bm{x}}(t)\in\mathbb{R}^{2d}:\dot{{\bm{x}}}=\{{\bm{x}},H\}\}$ denote a dataset of time-series samples, where each sample is a trajectory of the canonical coordinates ${\bm{x}}(t)$ evolving under Hamiltonian dynamics generated by an unknown Hamiltonian $H$. Throughout this work, we assume that the dataset $\bm{X}$} consists of simulated trajectories expressed in canonical coordinates. We prepare the dataset $\bm{X}$} to contain multiple trajectories initialized from diverse regions of phase space, so that the training procedure probes the global structure of the dynamics rather than a single dynamical regime.
The goal is to uncover the continuous Lie-group symmetries of such a classical dynamical system as defined by Eq. (5). To this end, we propose the Machine-Learning Symmetry Discovery (MLSD) algorithm, which uses neural networks (NNs) to parametrize each conserved quantity in Eq. (5) as $G_{i,\theta}$, where $\theta$ denotes the optimizable parameters of the neural networks.
@@
where $\epsilon$ denotes an infinitesimal variation and $\mathbf{J}$ is the symplectic metric defined in Eq. (4). Based on this principle, we can parametrize a conserved quantity $G_{i,\theta}$ and use its gradient, $\nabla_{{\bm{x}}}G_{i,\theta}$, to generate a symmetry transformation $\delta{\bm{x}}=\mathbf{J}(\nabla_{\bm{x}}G_{i})\epsilon$. In contrast to prior studies that directly model the symmetry transformation $\delta{\bm{x}}$ by neural networks[22, 14, 13, 23, 4], our approach of modeling $\delta{\bm{x}}$ indirectly through gradients of scalar functions $G_{i,\theta}$ is simpler, while inherently respecting the curl-free constraint $\nabla_{{\bm{x}}}\times(\nabla_{{\bm{x}}}G_{i,\theta})=0$ that should otherwise be imposed on $\delta{\bm{x}}$ as well. We remark that MLSD assumes access to trajectories expressed in canonical coordinates $(\mathbf{q},\mathbf{p})$, for which the symplectic form has the standard Darboux representation. Handling data of arbitrary coordinates would require learning the symplectic structure or a canonical transformation, which is beyond the scope of this work.
@@
    The symmetry discovery process begins by inputting a conjectured dimension $n$ of the Lie group. Therefore, $n+1$ individual neural networks are used to predict each conserved quantity. The structure coefficient $f^{ijk}$ is parametrized to be anti-symmetric, defined as $f^{ijk}=\eta^{ijk}-\eta^{ikj}+\eta^{jki}-\eta^{jik}+\eta^{kij}-\eta^{kji}$, for all $i,j,k\neq 0,$ with the parameter $\eta$ being optimizable. In order to recognize the symmetry group formed by all $G_{i,\theta}$, we train on the dataset $\bm{X}$} with the following mean square loss:
@@
    It should be understood that the integrand is evaluated instantaneously at each time $t$ and then integrated over the trajectory ${\bm{x}}(t)$ sampled from the dataset $\bm{X}$}.
In addition to $\mathcal{L}_{\text{H}}$ and $\mathcal{L}_{\text{G}}$, we introduce a third term in the loss function to encourage linear independence among the learned symmetry transformations. Without such a term, the optimization may converge to trivial solutions in which several $G_{i,\theta}$ correspond to equivalent conserved quantities. To quantify independence, we construct the $(n+1)\times 2d$ matrix
$\mathbf{M}({\bm{x}})=\big(\mathbf{J}\nabla_{{\bm{x}}}G_{0,\theta},\nabla_{{\bm{x}}}G_{1,\theta},\dots,\nabla_{{\bm{x}}}G_{n,\theta}\big)^{\intercal},$ (9)
and consider the eigenvalues $\mu^{i}({\bm{x}})\geq 0$ ($i=1,\dots,n+1$) of the Gram matrix $\mathbf{M}({\bm{x}})\mathbf{M}^{\intercal}({\bm{x}})$. We then define normalized eigenvalues
$\tilde{\lambda}^{i}({\bm{x}})=\frac{\mu^{i}({\bm{x}})}{\sum_{j=1}^{n+1}\mu^{j}({\bm{x}})},$ (10)
which satisfy $\tilde{\lambda}^{i}({\bm{x}})\in[0,1]$ and $\sum_{i=1}^{n+1}\tilde{\lambda}^{i}({\bm{x}})=1$. These normalized eigenvalues describe how the learned generators distribute their “weight” across independent directions in the space spanned by $\{\nabla_{{\bm{x}}}G_{i,\theta}\}$.
As an independence-promoting objective, we use the Shannon entropy of the normalized eigenvalues,
$\mathcal{S}_{\mathrm{ind}}({\bm{x}})=-\sum_{i=1}^{n+1}\tilde{\lambda}^{i}({\bm{x}})\log\!\big(\tilde{\lambda}^{i}({\bm{x}})+\varepsilon_{\lambda}\big),$ (11)
where a small constant $\varepsilon_{\lambda}$ is added inside the logarithm for numerical stability. In all experiments we set $\varepsilon_{\lambda}=10^{-5}$. Since the eigenvalues are normalized, this entropy is scale-invariant and remains well-behaved when some $\tilde{\lambda}^{i}$ approach zero. A higher value of $\mathcal{S}_{\mathrm{ind}}({\bm{x}})$ corresponds to a more isotropic eigenvalue spectrum, indicating that the learned generators span multiple independent directions, whereas a lower entropy signals collapse onto a lower-dimensional subspace. The corresponding independence term in the loss function is
$\mathcal{L}_{\text{I}}=-\mathop{\mathbb{E}}_{{\bm{x}}(t)\in{\bm{X}}}\int\mathcal{S}_{\mathrm{ind}}({\bm{x}}(t))\,\mathrm{d}t.$ (12)
In practice, $\mathcal{L}_{\text{I}}$ is weighted by a small factor $\beta\ll 1$ in the total loss, so that it acts as a weak regularizer that discourages degeneracy among generators without dominating the Hamiltonian-learning and algebra-closure objectives. Similar entropy-based independence regularization has also been explored in [12].
In summary, the overall loss function is
@@
Throughout this work we set $\alpha=1$, so that $\mathcal{L}_{\text{H}}$ and $\mathcal{L}_{\text{G}}$ contribute at comparable magnitude, while the independence term $\mathcal{L}_{\text{I}}$ is weakly weighted with $\beta=10^{-4}$.
After training, if the dataset $\bm{X}$} corresponds to a system with a continuous $n$-dimensional symmetry, the first two loss terms, Eq. (7) and Eq. (8), are expected to approach zero, while the independence term Eq. (12) remains finite. Conversely, if the system does not possess an $n$-dimensional symmetry, the first two loss terms remain bounded away from zero.
In practice, even when the correct symmetry dimension is supplied, $\mathcal{L}_{\text{H}}$ and $\mathcal{L}_{\text{G}}$ may not converge exactly to zero due to the finite expressive capacity of the neural networks and optimization limitations. Nevertheless, a clear separation in the converged loss values is observed between the true symmetry dimension and mismatched (over- or under-parameterized) cases. The specific symmetry group is then identified by extracting the structure coefficients $f^{ijk}$, as detailed in Appendix B.
@@
Harmonic oscillator $H=\frac{{\bm{p}}^{2}}{2}+\frac{{\bm{q}}^{2}}{2}$ $G_{i}=\frac{1}{2}({\bm{q}}+\mathrm{i}{\bm{p}})^{T}M_{i}({\bm{q}}-\mathrm{i}{\bm{p}}),$ $M_{i}\in\text{Gell-Mann matrices}.$ $\{G_{i},G_{j}\}=\frac{1}{2}({\bm{q}}+\mathrm{i}{\bm{p}})^{T}[M_{i},M_{j}]({\bm{q}}-\mathrm{i}{\bm{p}}).$ SU(3)
@@ III Empirical results
In this section, we apply MLSD to discover symmetries in the harmonic oscillator and the Kepler problem, summarized in Tab. 1. A detailed analysis of the $\mathrm{SU}(3)$ symmetry of the harmonic oscillator and the $\mathrm{SO}(4)$ symmetry of the Kepler problem is provided in Appendix A. We demonstrate that MLSD can successfully identify the hidden symmetries in both systems[6]. Additionally, in the harmonic oscillator task, we use a quadratic parametrization to show that MLSD can explicitly reconstruct the eight Gell–Mann matrices that form the $\mathfrak{su}(3)$ Lie algebra.
Throughout this section, all trajectory rollouts are generated using a symplectic integration scheme. In our implementation, both the ground-truth Hamiltonians and the learned Hamiltonian $G_{0,\theta}$ are parameterized in a separable form,
$H(\mathbf{q},\mathbf{p})=T(\mathbf{p})+V(\mathbf{q}),$ (14)
which enables the use of the velocity-Verlet (leapfrog) symplectic integrator. The integration scheme advances the canonical variables as
$\mathbf{p}_{n+\frac{1}{2}}$ $=\mathbf{p}_{n}-\frac{h}{2}\nabla_{\mathbf{q}}V(\mathbf{q}_{n}),$ (15)
$\mathbf{q}_{n+1}$ $=\mathbf{q}_{n}+h\,\nabla_{\mathbf{p}}T(\mathbf{p}_{n+\frac{1}{2}}),$
$\mathbf{p}_{n+1}$ $=\mathbf{p}_{n+\frac{1}{2}}-\frac{h}{2}\nabla_{\mathbf{q}}V(\mathbf{q}_{n+1}).$
where $h$ denotes the integration time step. This integrator is symplectic by construction, as it corresponds to a composition of exact Hamiltonian flows generated separately by $T(\mathbf{p})$ and $V(\mathbf{q})$, and ensures reliable long-time energy conservation for both the learned and ground-truth dynamics. Energy conservation is assessed by tracking the deviation
$\Delta H(t_{n})\equiv\big|H(\mathbf{q}_{n},\mathbf{p}_{n})-H(\mathbf{q}_{0},\mathbf{p}_{0})\big|,$ (16)
along each trajectory, which provides a direct measure of energy drift in the learned dynamics.
To quantify energy conservation beyond time-resolved drift curves, we also compute the root-mean-square (RMS) energy deviation along each trajectory,
$\Delta H_{\mathrm{RMS}}=\sqrt{\frac{1}{N}\sum_{n=1}^{N}\big[H(\mathbf{q}_{n},\mathbf{p}_{n})-H(\mathbf{q}_{0},\mathbf{p}_{0})\big]^{2}},$ (17)
where $N$ denotes the number of integration steps. The RMS deviation provides a single scalar measure that summarizes the typical magnitude of energy drift accumulated over the trajectory. For clarity, the RMS energy deviation for each trajectory is reported as an inset in the corresponding energy-drift panels.